On June 30, 2021, Georgia-Atlantic, Inc. leased warehouse equipment from Builders, Inc. The lease agreement calls for Georgia-Atlantic to make semiannual lease payments of $880,440 over a 3-year lease term (also the asset’s useful life), payable each June 30 and December 31, with the first payment at June 30, 2021. Georgia-Atlantic's incremental borrowing rate is 8.0%, the same rate Builders used to calculate lease payment amounts. Builders manufactured the equipment at a cost of $4.3 million. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1. Determine the price at which Builders is “selling” the equipment (present value of the lease payments) at June 30, 2021.
2. What amount related to the lease would Builders report in its balance sheet at December 31, 2021 (ignore taxes)?
3. What line item amounts related to the lease would Builders report in its income statement for the year ended December 31, 2021 (ignore taxes)?
(For all requirements, enter your answers in whole dollars and not in millions. Round your final answer to the nearest whole dollar.)
Answer #1

1. Determine the price at which Builders is “selling” the equipment (present value of the lease payments) at June 30, 2021.

$ 4,800,002.45 = (NPV(8%/2,880440,880440,880440,880440,880440)+880440


2. What amount related to the lease would Builders report in its balance sheet at December 31, 2021 (ignore taxes)?

$ 3,195,904.95 = (4,800,002.45 -880440- (880440-(4,800,002.45-880440)*4%))


3. What line item amounts related to the lease would Builders report in its income statement for the year ended December 31, 2021 (ignore taxes)?

Revenue : 4,800,002.45

Cost of Good Sold : 4,300,000

Interest income from Lease = 156,782.50 (4,800,002.45-880440)*4%)